# Observation of non-Hermitian Dirac cones

### v1

@@
In summary, we have observed non-Hermitian Dirac cones in a graphene-like model with nonreciprocal NNN couplings. Protected by the TRS, Dirac quasiparticles in different valleys acquire highly contrasting lifetimes, leading to intriguing non-Hermitian valley contrasting physics. Using circuiting experiments, we demonstrate the effective single Dirac cone behavior of our model, which allows for auto valley selection and vortex state generation. At the boundary, topological valley kink states are found to inherit the non-Hermitian physics in the bulk and exhibit unidirectional propagation and suppressed inter-valley scattering that are not seen in the Hermitian case. These results demonstrate that Dirac quasiparticles can not only exist in realistic non-Hermitian settings but also have profound physical implications.
@@
Our work takes the first step in studying non-Hermitian quasiparticles and extends the current studies on non-Hermitian topological semimetal phases which are mostly focused on exceptional points and conventional Hermitian band degeneracies. There are many efforts to be made along this direction. At the fundamental level, it would be interesting to search for other types of non-Hermitian quasiparticles, especially non-Hermitian Weyl fermions, which have been predicted to host exotic three-dimensional non-Hermitian physics [37, 38, 22, 39, 40]. Practically, it is highly desired to extend our results to wave (e.g., microwave, optical and acoustic) and even electronic systems, where various valley-based applications may be further developed.
@@ Methods
The data that support the findings of this study are available from the corresponding authors upon reasonable request.
@@
The codes that support the findings of this study are available from the corresponding authors upon reasonable request.
@@ Acknowledgements
We are grateful to Yidong Chong and Baile Zhang for fruitful discussions. X.X., F.M., Z.D., Y.D., H.C. and F.G. are supported by the Key Research and Development Program of the Ministry of Science and Technology under grants No. 2022YFA1404902, 2022YFA1404704 and 2022YFA1405200, the National Natural Science Foundation of China (NNSFC) under grants No. 62171406, 11961141010 and 61975176, the Zhejiang Provincial Natural Science Foundation under grant No. Z20F010018, the Key Research and Development Program of Zhejiang Province under grant No. 2022C01036, and the Fundamental Research Funds for the Central Universities. H.X. acknowledges the support of the start-up fund and the direct grant (grant No. 4053675) from The Chinese University of Hong Kong. W.B.R. is supported by the RGC Postdoctoral Fellowship (Ref. No. PDFS2223-7S05).

### v2

@@ Discussion
In summary, we have observed non-Hermitian Dirac cones in a graphene-like model with nonreciprocal NNN couplings. Protected by the TRS, Dirac quasiparticles in different valleys acquire highly contrasting lifetimes, leading to intriguing non-Hermitian valley contrasting physics. Using circuiting experiments, we demonstrate the effective single Dirac cone behavior of our model, which allows for auto valley selection and vortex state generation. At the boundary, topological valley kink states are found to inherit the non-Hermitian physics in the bulk and exhibit unidirectional propagation and suppressed inter-valley scattering that are not seen in the Hermitian case, which could be used for robust logic gates and compact circuit integration. These results demonstrate that Dirac quasiparticles can not only exist in realistic non-Hermitian settings but also have profound physical implications.
@@
Our work takes the first step in studying non-Hermitian quasiparticles and extends the current studies on non-Hermitian topological semimetal phases which are mostly focused on exceptional points and conventional Hermitian band degeneracies. There are many efforts to be made along this direction. At the fundamental level, it would be interesting to search for other types of non-Hermitian quasiparticles and their interactions with external fields [38, 39], especially non-Hermitian Weyl fermions, which have been predicted to host exotic three-dimensional non-Hermitian physics [40, 41, 22, 42, 43]. Practically, it is highly desired to extend our results to wave (e.g., microwave, optical and acoustic) and even electronic systems, where various valley-based applications may be further developed.
@@ Methods
All the data supporting this study are available in the paper and Supplementary Information. Additional data related to this paper are available from the corresponding authors upon request.
@@
The codes that support the findings of this study are available from the corresponding authors upon request.
@@ Acknowledgements
We are grateful to Yidong Chong and Baile Zhang for fruitful discussions. X.X., F.M., Z.D., Y.D., H.C. and F.G. are supported by the Key Research and Development Program of the Ministry of Science and Technology under grants No. 2022YFA1404902, 2022YFA1404704 and 2022YFA1405200, the National Natural Science Foundation of China (NNSFC) under grants No. 62171406, 11961141010 and 61975176, the Zhejiang Provincial Natural Science Foundation under grant No. R25F010016, the Key Research and Development Program of Zhejiang Province under grant No. 2022C01036, and the Fundamental Research Funds for the Central Universities. H.X. acknowledges the support of the start-up fund and the direct grant (grant No. 4053675) from The Chinese University of Hong Kong. W.B.R. is supported by the RGC Postdoctoral Fellowship (Ref. No. PDFS2223-7S05).